OpenStudy (anonymous):

Of all the people applying for a certain job, 90% are qualified and 10% are not. The personnel manager claims that she approves qualified people 80% of the time, she approves unqualified people 40% of the time. Find the probability that a person is qualified if he or she was not approved by the manager.

OpenStudy (kropot72):

\[P(qualified\ but\ unapproved)=0.9\times 0.2\] \[P(unqualified\ but\ unapproved)=0.1\times 0.6\] Can you find the probability of a qualified or an unqualified person being unapproved?

OpenStudy (anonymous):

P(qualified but unapproved)= 0.18 P(unqualified but unapproved)= 0.06 can u show me what i need to do next?

OpenStudy (kropot72):

P(qualified or unqualified person being not approved) = 0.18 + 0.06 = 0.24 So the probability of any applicant being not approved = 0.24 and the probability of a qualified applicant being not approved is 0.18. Therefore the probability that a person is qualified if he or she was not approved by the manager is 0.18/0.24 = ?

OpenStudy (anonymous):

0.75 :)

OpenStudy (kropot72):

Looks good to me :)